The math club is selling pizza and lemonade to raise money for a field trip. the trip will cost $450
If they sell pizza for $1 per slice for and lemonade$2 a bottle and sold 190 items. How many of each did they sell?
The sum of 2 numbers is 87. the first number is 3 more than twice the second number.
2x+3y=12
X=y+4

Answer by TimothyLamb(4379)

---
you have 3 different problems here:
---
first problem (pizza and lemonade):
---
x + 2y = 450
x + y = 190
y = 190 - x
---
x + 2(190 - x) = 450
x + 380 - 2x = 450
---
x = pizza slices sold = -70 (nobody can sell a negative number of pizza slices)
y = lemonades sold = 260
---
answer:
the problem statement contains an error because the equations yield a nonsensical solution.
---
second problem:
x + y = 87
x = 2y + 3
---
put the system of linear equations into standard form
---
x + y = 87
x - 2y = 3
---
solve by matrix method (the below solver uses the matrix method):
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
answer:
x= 59
y= 28
---
solve by elimination:
---
x + y = 87
-x + 2y = -3
---
add:
3y = 84
---
answer (same as obtained by the matrix method above):
y = 28
x = 59
---
third problem:
---
2x + 3y = 12
x = y + 4
---
put the system of linear equations into standard form
---
2x + 3y = 12
x - y = 4
---
solve by matrix method (the below solver uses the matrix method):
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
answer:
x= 4.8
y= 0.8
---
solve by elimination:
---
2x + 3y = 12
-2x + 2y = -8
---
add:
5y = 4
---
answer (same as obtained by the matrix method above):
y = 4/5 = 0.8
x = 4.8
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php